OpenStudy (anonymous):

(2a^3b^-2) (-4a^2b^4) i dont understand..

OpenStudy (anonymous):

step by step what is \(2\times -4\) ?

OpenStudy (anonymous):

−8a^5b^2

OpenStudy (anonymous):

could you explain?

OpenStudy (anonymous):

\(2\times -4=-8\)\[a^2\times a^3=a^{2+5}=a^5\] \[b^4\times b^{-2}=b^{4-2}=b^2\]

OpenStudy (anonymous):

i thought i had to get rid of negative powers first

OpenStudy (anonymous):

nah, just add up the exponents for the like terms

OpenStudy (anonymous):

soo 2*-4 then add the exponents n then 1b+b1 then add-2 and 4 ?

OpenStudy (anonymous):

not sure what you mean by 1b+b1

OpenStudy (anonymous):

dont the bs have an understood one

OpenStudy (anonymous):

\[\large (2a^3b^{-2}) (-4a^2b^4)\] lets put the like terms next to each other \[\large 2\times -4\times a^3\times a^2\times b^{-2}\times b^4\]

OpenStudy (anonymous):

clear where the \(-8\) comes from right?

OpenStudy (anonymous):

2a^3+-4a^2

OpenStudy (anonymous):

oh i see the confusion it is all multiplication, there is not addition here

OpenStudy (anonymous):

simplifing

OpenStudy (anonymous):

the coefficients (numbers) are not attached to the variables. you multiply them together, they are numbers

OpenStudy (anonymous):

it is a multiplication, not an addition btw there is no such mathematical operation as "simplify" that is the lazy math teachers way of writing "give me an answer i want" you can add, subtract, multiply, divide in this case you are supposed to multiply

OpenStudy (anonymous):

but here when you multiply \(a^3\times a^3\) you get \(a^{2+3}=a^5\) i.e. you add the exponents

OpenStudy (anonymous):

oops typo there, i meant \(a^2\times a^2=a^{2+3}=a^5\)

OpenStudy (anonymous):

oops my bad i forgot :/

OpenStudy (anonymous):

\[\overbrace{a\times a}^{a^2}\times \overbrace{a\times a\times a}^{a^3}=\overbrace{a\times a\times a\times a\times a}^{a^5}\]

OpenStudy (anonymous):

i dont understand that

OpenStudy (anonymous):

the meaning of \(a^2\) is \(a\times a\) right?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

similarly \(a^3\) means \(a\times a\times a\)

OpenStudy (anonymous):

oh ok ya i did know that

OpenStudy (anonymous):

that was my explanation of why \(a^2\times a^3=a^5\)

OpenStudy (anonymous):

similarly \(b^{-2}\times b^4=b^{-2+4}=b^2\)

OpenStudy (anonymous):

so for this problem you had to do three things: \(2\times -4=-8\) and \(a^2\times a^3=a^5\) and also \(b^{-2}\times b^4=b^2\) giving a final answer of \[-8a^5b^2\]

OpenStudy (anonymous):

ohh ok gotcha thankyou